OpenStudy (anonymous):

A tennis court is 36 feet wide. Its length is 6 feet more than twice the width. What is the area in square feet of a tennis court?

OpenStudy (anonymous):

your given width = 36 length = 6 + 2w = 6+2(36) = 78 area = length(width) area = 36(78) = 2808sq ft

OpenStudy (anonymous):

L=36*2=6 so L*W equals __________